3. The spectrum shown here represents one of the compounds (A-E) from question 2. (a) Which compound? (b) Explain your reasoning. 4. Visit an online spectral database (SDBS, NIST, etc.) and find the IR spectrum for 2-methylbutanal. Insert the spectrum here.
Strang Subtitie Title No Spacing For each structure A-E, determine whether an absorption for the functional groups shown in the table 2. will be present in its infrared spectrum, If not, leave the box blank. If yes, fill in the appropriate boxes with the expected frequencies (in cm). Structure C-C Aldehyde Ketone Alcohol Aromatic (not aromatie) A 1680-1660 KC. CH B 1720 C 1640 CHy OCH D 860 1640-1640 1690 860 1630-1610
Answer #1

3 a) the given IR spectrum represents compound D (1-methoxy-4-((E)-prop-1-enyl)benzene)

ОСНЗ CНз

b) reason : the major bands in the spectrum confirms the structure of 1-methoxy-4-((E)-prop-1-enyl)benzene. They are: the bands at 1460–1609 cm-1 is due to aromatic C=C stretching. The aromatic –C-H appears as a band in 3005 cm-1. –CH3 stretching is observed at 2850-2920 cm-1. Non-aromatic C=C appears at 1653 cm-1. The aromatic ether (-C-O-C) part is observed at 1243 cm-1.

4) IR spectrum of 2-methylbutanal is given below

LOD pрon 5D 3000 е00л 1000 15 ЛО 4роо 5 о НАVENUHB ERI-I

2)

A) Aromatic (1400-1600 cm-1 for stretch and 1500-1700 cm-1), C=C non-aromatic (1640-1680 cm-1)

B) Aldehyde (1720-1740 cm-1), Aromatic (1400-1600 cm-1 for stretch and 1500-1700 cm-1).

C) Alcohol (O-H stretch 3200-3600 cm-1, C-O stretch 1050-1150 cm-1)

Ether (C-O stretch 1050-1150 cm-1)

C=C non-aromatic (1640-1680 cm-1)

D) Ether (C-O stretch 1050-1150 cm-1)

Aromatic (1400-1600 cm-1 for stretch and 1500-1700 cm-1), C=C non-aromatic (1640-1680 cm-1)

E) ketone (C-O stretch 1680-1750 cm-1)